# Selective Prompt Anchoring for Code Generation

### v5

@@ 5.6 Impact of Anchoring Strength
In this section, we study how the anchoring strength of Spa affects performance. To ensure fair comparisons, Spa is activated for all tasks in these experiments, unlike during inference when it is only activated for failed tasks. We also tune Spa on all tasks to improve generalizability. As shown in Figure 3, represents the original model.
We find that different anchoring strengths have different effects: if is too low, improvements are low; if too high, the LLM starts becoming biased and performance declines. Overall, as increases, performance first improves, reaches an optimum, and then drops with further increases.
We find is both tunable and transferable. While the optimal value varies slightly across models and benchmarks, it is generally model-dependent. We find serves as a universal anchoring strength, consistently improving performance in all cases. Additional experiments and discussions are provided in Appendix J.
@@ Acknowledgements
A Performance Analysis by Different Generation Lengths A
B Calculation Details and Extended Discussion of LLM Attention B
B.1 Calculation of Self-attention B.1
B.2 Calculation of Gradient-based Attention B.2
B.3 Attention Misalignment B.3
B.4 Attention to the User Prompt Ratio B.4
C Extended Discussion of Approximation in Spa C
D Code Generation Examples D
E Computational Cost of Spa E
E.1 Latency E.1
E.2 Memory E.2
E.3 Tradeoff Between Latency and Memory Usage via Parallelization E.3
F Pass@10 & Beam Search with Spa F
G Hypothetical explanation for Attention dilution and SPA’s effectiveness G
H Implementation and Deployment H
H.1 Imlementation of Spa H.1
H.2 Model Deployment H.2
H.3 Prompt Design H.3
I Evaluating Spa on Other Generative Tasks I
J Tuning of Anchoring Strength J
J.1 Tuning Stability J.1
J.2 Cross-dataset & Cross-Model Tuning Evaluation J.2
J.3 Comparision between Tuned Spa and Optimal Spa J.3
K Analysis of Spa’s Performance without Test Case Availability K
L Comparison to Pasta L

### v6

@@ 5.6 Impact of Anchoring Strength
In this section, we discuss how the anchoring strength of Spa influences performance. We observe that different values of yield different effects: when is too low, performance gains are limited; when it is too high, the LLM becomes biased, leading to performance degradation. As shown in Figure 3, corresponds to the original generation. As increases, performance initially improves, reaches an optimum, and subsequently declines with further increases in . This simple pattern applies to all the settings. We find is both tunable and transferable. While the optimal value may vary slightly across models and benchmarks, it is generally model-dependent. A value of consistently yields robust improvements across all settings. Further details are provided in Appendix J.
To make Spa more stable and eliminate the need of tuning, we also propose a confidence-modulated weighting strategy that adjusts each token’s anchoring strength based on its original probability. Details are provided in Appendix M.
@@ Appendix I Evaluating Spa on Other Generative Tasks
Appendix M Confidence-Modulated Anchoring Strength
To make Spa more stable and eliminate the need for tuning, we propose a confidence-modulated weighting method that adjusts each token’s anchoring strength based on its original probability respectively.
When the LLM is highly confident about a specific token, the corresponding logit is adjusted less, resulting in a lower anchoring weight applied to the logit difference. Conversely, if the LLM shows lower confidence, a relatively higher weight is applied, leading to a greater adjustment of the logit value. The confidence level is measured using the original probability distribution generated by the LLM.
Formally, let represent the predicted probability of token within the original probability distribution for the next token. The anchoring strength for token is defined as:
[…]